im trying to do the elimination method for this problem: 2x+6y=8
3x+9y=12
please help thanks!

Answer by ReadingBoosters(3246)

2x+6y=8
3x+9y=12
...
Multiply the first equation by 3 and the second by -2, to get
3(2x+6y=8)
-2(3x+9y=12)
...
6x + 18y = 24
-6x - 18y = -24
...
Add the equatins
6x + -6x + 18y + -18y = 24 + -24
0 = 0
...
Meaning that these are the same line, intersecting at every point.
It is a dependent system.
...
CHECK
2x + 6y = 8
6y = -2x + 8
y = (-2/6)x + (8/6)
y = (-1/3)x + (4/3)-->linear equation
...
3x+9y=12
9y = -3x + 12
y = (-3/9)x + 12/9
y (-1/3)x + 4/3 --> same linear equation
Hence, they are the same line, where the solution is the entire line.
.....................
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.